# RACOON: An LLM-based Framework for Retrieval-Augmented Column Type Annotation with a Knowledge Graph

### v1

@@ Abstract
As an important component of data exploration and integration, Column Type Annotation (CTA) aims to label columns of a table with one or more semantic types. With the recent development of Large Language Models (LLMs), researchers have started to explore the possibility of using LLMs for CTA, leveraging their strong zero-shot capabilities. In this paper, we build on this promising work and improve on LLM-based methods for CTA by showing how to use a Knowledge Graph (KG) to augment the context information provided to the LLM. Our approach, called Racoon, combines both pre-trained parametric and non-parametric knowledge during generation to improve LLMs’ performance on CTA. Our experiments show that Racoon achieves up to a 0.21 micro F-1 improvement compared against vanilla LLM inference.
@@ 1 Introduction
In recent years, researchers and data scientists have found abundant relational tables on the Web (Cafarella et al. (2008a, b); Bhagavatula et al. (2015)). Despite their high quality, these Web Tables often miss important meta information such as column headers and relationships between columns (Suhara et al. (2022)). This information is essential for downstream tasks such as data quality control (Schelter et al. (2018)) and data discovery (Chapman et al. (2020)). A key step in recovering this meta information is Column Type Annotation (CTA), which aims to assign one or more semantic type(s) to columns in a given table.
@@
There has been much research effort in developing effective solutions that utilize LLMs for CTA and related tasks. One initial attempt to apply LLMs for various table-related tasks is from (Narayan et al. (2022)). For CTA specifically, one line of work (Korini and Bizer (2023)) directly applies LLMs to solve CTA with promising performance. CHORUS (Kayali et al. (2024)) and Archetype (Feuer et al. (2023)) further develop other components using history and remapping to improve performances. Table-GPT (Li et al. (2024)) fine-tunes LLMs to adapt to relational data format to improve performance on various table understanding tasks.
@@
Despite promising performance, LLM-based methods still encounter challenges dealing with outdated knowledge (He et al. (2022)), producing factual inaccuracies (Ji et al. (2023)), and handling domain-specific or specialized queries (Kandpal et al. (2023)). To alleviate such problems, Retrieval-Augmented Generation (RAG), which enhances content generated by LLMs by retrieving knowledge from external sources as non-parametric memory (Asai et al. (2023)), has emerged as a promising approach. Among the many external knowledge sources, Knowledge Graphs (KGs), can provide succinct yet informative knowledge to help analyze the intricate semantics of entities (Ji et al. (2021); Wang et al. (2017)). Specifically, KGs represent information in the form of entities (nodes) and relations (edges) (Wang et al. (2017)). Each edge is a factual triplet in the form of (subject, relation, object), making KG a well-structured data source. Many recent approaches have studied integrating KGs with LLMs/PLMs through RAG to improve performances on various tasks (Xu et al. (2020); Pan et al. (2024); Dehghan et al. (2024); Sun et al. (2024); Wen et al. (2024)) and saw promising performance gains compared with vanilla LLM inference. But none of the above approaches focuses on CTA. The most relevant work KGLink ( Wang et al. (2024)) incorporates an external KG with PLM during training for CTA, thus encountering the exact PLM problems we mentioned before, whereas our work aims to leverage LLMs and KG to avoid expensive training or fine-tuning over large models.
Figure 1: Vanilla way of using LLM for CTA (left) vs Racoon (right). Racoon processes the input table, and based on the cell information, it retrieves from an external knowledge graph necessary knowledge as additional context information to augment the prompt for the LLM.
@@
Inspired by these advances, we argue that we can similarly bring external non-parametric knowledge to LLMs for CTA to improve performance and propose our framework Racoon. As shown in Figure 1, compared with vanilla LLM inference on the left, Racoon on the right unifies LLM inference with a KG by performing three additional steps: namely after ingesting a table as the input, Racoon first (1) looks at the column cells’ entity mention and retrieves related information from a KG. However, this retrieved content cannot be directly applied due to its large size and noise. Thus Racoon needs to (2) process the retrieved content, and finally (3) serialize the compressed context to augment the prompt for the LLM.
@@
(1) We introduce the problem of augmenting prompts for LLMs on CTA with external knowledge from a KG and propose an end-to-end framework Racoon to solve this problem.
(2) We explore different granularity of information Racoon can retrieve from a KG for RAG. Furthermore, we provide an effective post-retrieval compression method and a serialization method such that given a column of cells, we can derive additional context information to augment the prompt.
@@ 3 Experiments
KG-Linker Baseline Ent. Label Ent. Triplets Baseline Ent. Label Ent. Triplets
Ground Truth 0.2609 0.4023 0.45 0.4631 0.564 0.6814
MediaWiki API 0.2609 0.318 0.284 0.4631 0.4886 0.4336
ReFinED 0.2609 0.3698 0.3678 0.4631 0.49 0.5705
Evaluation. We first use the ground truth entity ID linked to each cell labeled in the WikiTables-TURL-CTA dataset as the output of the KG-Linker to retrieve information from the KG. This enables us to evaluate our approach assuming a perfect KG-Linker. The results show that both augmented prompts using Entity-Triplets and Entity-Labels outperform vanilla LLM inference, with Entity-Triplets achieving the best performance overall. This highlights the potential of retrieving more comprehensive information about column cells from the KG.

### v2

@@ Abstract
As an important component of data exploration and integration, Column Type Annotation (CTA) aims to label columns of a table with one or more semantic types. With the recent development of Large Language Models (LLMs), researchers have started to explore the possibility of using LLMs for CTA, leveraging their strong zero-shot capabilities. In this paper, we improve on LLM-based methods for CTA by showing how to use a Knowledge Graph (KG) to augment the context information provided to the LLM. Our approach, called Racoon, combines both pre-trained parametric and non-parametric knowledge during generation to improve LLMs’ performance on CTA. Our experiments show that Racoon achieves up to a 0.21 micro F-1 improvement compared against vanilla LLM inference.
@@ 1 Introduction
In recent years, researchers and data scientists have found abundant relational tables on the Web (Cafarella et al. (2008a, b); Bhagavatula et al. (2015)). Despite their high quality, these Web Tables often miss important meta information such as column headers and relationships between columns (Suhara et al. (2022)). This information is essential for downstream tasks such as data quality control (Schelter et al. (2018)) and data discovery (Chapman et al. (2020)). A key step in recovering this meta information is Column Type Annotation (CTA), which aims to assign one or more semantic type(s) to columns in a given table. For example, the column [’UEFA Champions League’, ’Scottish Cup’] has the annotation “time.event”.
@@
There has been much research effort in developing effective solutions that utilize LLMs for CTA and table related tasks (Narayan et al. (2022), Tian et al. (2024)). For CTA specifically, one line of work (Korini and Bizer (2023)) directly applies LLMs to solve CTA with promising performance. CHORUS (Kayali et al. (2024)) and Archetype (Feuer et al. (2023)) further develop other components using history and remapping to improve performances. Table-GPT (Li et al. (2024)) fine-tunes LLMs to adapt to relational data format to improve performance on various table understanding tasks.
@@
At the core of existing methods, they directly apply LLMs without any external knowledge. Despite promising performance, LLM-based methods still encounter challenges dealing with outdated knowledge (He et al. (2022)), producing factual inaccuracies (Ji et al. (2023)), and handling domain-specific or specialized queries (Kandpal et al. (2023)). To alleviate such problems, Retrieval-Augmented Generation (RAG), which enhances LLMs’ generation by retrieving knowledge from external sources as non-parametric memory (Asai et al. (2023)), has emerged as a promising approach. Among the many external knowledge sources, Knowledge Graphs (KGs), can provide succinct yet informative knowledge to help analyze the intricate semantics of entities (Ji et al. (2021); Wang et al. (2017)). Specifically, KGs represent information in the form of entities (nodes) and relations (edges) (Wang et al. (2017)). Each edge is a factual triplet in the form of (subject, relation, object), making KG a well-structured data source. Many recent approaches have studied integrating KGs with LLMs/PLMs through RAG to improve performances on various tasks (Xu et al. (2020); Pan et al. (2024); Dehghan et al. (2024); Sun et al. (2024); Wen et al. (2024)) and saw promising performance gains compared with vanilla LLM inference. But none of the above approaches focuses on CTA. The most relevant work KGLink ( Wang et al. (2024)) incorporates an external KG with PLM during training for CTA, thus encountering the exact PLM problems we mentioned before, whereas our work aims to leverage LLMs and KG to avoid expensive training or fine-tuning over large models.
Figure 1: Vanilla way of using LLM for CTA (left) vs Racoon (right).
@@
Inspired by these advances, we argue that we can similarly bring external non-parametric knowledge to LLMs for CTA to improve performance and propose our framework Racoon. As shown in Figure 1, compared with vanilla LLM inference on the left, Racoon on the right unifies LLM inference with a KG by performing three additional steps: namely after ingesting a table as the input, Racoon first (1) looks at the column cells’ entity mention and retrieves related information from a KG. However, this retrieved content cannot be directly applied due to its large size and noise. Thus Racoon needs to (2) process the retrieved content, and finally (3) serialize the compressed context to augment the prompt for the LLM. Note that Racoon is orthogonal to existing LLM-based methods for CTA, which enhance LLMs’ performance through additional demonstrations and response post-processing.
@@
(1) We introduce the problem of augmenting prompts for LLMs on CTA with external knowledge from a KG and propose an end-to-end framework Racoon for it.
(2) We explore different granularity of information Racoon can retrieve from a KG with RAG. Furthermore, we provide effective post-retrieval compression and serialization methods such that given a column of cells, we can derive additional context information to augment the prompt.
@@ 3 Experiments
KG-Linker EL Baseline ETL ETT Baseline ETL ETT
Ground Truth 1.0000 0.2609 0.4023 0.4500 0.4631 0.564 0.6814
MediaWiki API 0.4018 0.2609 0.3582 0.3494 0.4631 0.4927 0.5222
ReFinED 0.6255 0.2609 0.3698 0.3678 0.4631 0.4900 0.5705
Evaluation. We first use the ground truth entity ID linked to each cell labeled in the WikiTables-TURL-CTA dataset as the output of the KG-Linker to retrieve information from the KG. This enables us to evaluate our approach assuming a perfect KG-Linker. The results in Table 1 show that both augmented prompts using Entity-Labels (ETL) and Entity-Triplets (ETT) outperform vanilla LLM inference, with Entity-Triplets achieving the best performance overall. This highlights the potential of retrieving more comprehensive information about column cells from the KG.
@@ 4 Conclusion
Acknowledgments
This project was funded in part by a gift from NEC.
Appendix A Appendix
Here we show the example prompts for the single-label setting and the multi-label setting.
A.1 Single-label
System message: Be a helpful, accurate assistant for data discovery and exploration desiged to output valid JSON in the format {‘type’: []}
User message: Consider this table given in Comma-separated Values format: “‘ Table “‘ There are a list of 255 valid types for each column: types. Your task is to choose only one type from the list to annotate the first column. Solve this task by following these steps: 1. Look at the cells in the first column of the above table. 2. Consider this information carefully: Cells in this column are instances of the following wikidata entities: human (6 cells). 3. Choose only one valid type from the given list of types. Check that the type MUST be in the list. Give the answer in valid JSON format.
A.2 Multi-label:
System message: Be a helpful, accurate assistant for data discovery and exploration desiged to output valid JSON in the format {‘type’: []}
User message: For multi-label: Consider this table given in Comma-separated Values format: “‘ Table “‘ There are a list of 255 valid types for each column: types. Your task is to choose one or multiple types from the list to annotate the first column. Solve this task by following these steps: 1. Look at the cells in the first column of the above table. 2. Consider this information carefully: Cells in this column are instances of the following wikidata entities: human (6 cells). 3. Mark each type in the given list with 0 or 1. Mark a type with 1 if it can better represent all cells of the first column. Mark a type with 0 otherwise. 4. Give a list of types that you have marked 1 in the previous step. Check that the types MUST be in the list. Give the answer in valid JSON format.